Systems and Network Management

The Structure of Management Information (SMI)

1 Background

The snmp protocol is called “simple” because the protocol itself is quite simple. However, the difficulty is in applying it to actually managing systems and networks.

There are many terms and standards involved; it is necessary to understand enough of them to make sense of the mibs that define the objects that you want to monitor and manage. If you can make sense of the mib files, you can identify the objects that you want to monitor.

1.1 Management Information Base (MIB)

The mibs define the objects that you can manage.

When you installed the Net snmp software package, you installed some mib files into the directory /usr /share/snmp/mibs/. You can list them all with:

$ rpm -ql ucd-snmp | grep snmp/mibs/.*\.txt

There are many other mibs that are not included here; you can download others from somewhere such

as http://www.simpleweb.org/ietf/ and include
them into your Net snmp clients as explained at http://net-snmp.sourceforge.net/FAQ.html#How_do_I_add_a_MIB_
and at http://net-snmp.sourceforge.net/tutorial/commands/mib-options.html.

Nick Urbanik <nicku(at)nicku.org>  ver. 1.1

The Structure of Management Information (SMI) Systems and Network Management

2

1.2 Structure of Management Information

smi is a definition of the structure of the mibs, how they are connected together into a tree. See the rfcs below in section 2.1 on page 10. It specifies which part of asn.1 will be used to define mibs.

1.3 Abstract Syntax Notation One (ASN.1)

asn.1 is widely used for many things other than snmp.

See http://asn1.elibel.tm.fr/en/uses/ for a list

of some of the applications of asn.1. There is a web site dedicated to providing information about it at

http://asn1.elibel.tm.fr/.

1.4 Basic Encoding Rules (BER)

The basic encoding rules is an iso standard. It describes a method for encoding values of each asn.1 type as a string of octets.

1.5 What we are doing today

We will examine the specification for mib-2, on your

machine at /usr/share/snmp/mibs/RFC1213-MIB.txt

and understand the structure of it.

1.6 Syntax of a Managed Object Definition

Every object definition in smi has the format:

name OBJECT-TYPE
    SYNTAX datatype

Nick Urbanik <nicku(at)nicku.org>  ver. 1.1

The Structure of Management Information (SMI) Systems and Network Management

3

    ACCESS either read-only, read-write, write-only. or not-accessible
    DESCRIPTION
        "Some text that describes this managed object."
    ::= { unique object ID that defines this object }

We will refer to this later in our activities.

2 The MIB-II Definition

Here I will refer to my edited version of RFC1213-MIB.txt,

available at http://nicku.org/snm/lectures/smi/RFC1213-MIB.txt.
RFC1213-MIB DEFINITIONS ::= BEGIN
IMPORTS
        mgmt, NetworkAddress, IpAddress, Counter, Gauge,
                TimeTicks
            FROM RFC1155-SMI
        OBJECT-TYPE
                FROM RFC-1212;

The first line defines the name of the mib, here RFC1213-MI. The format of this definition is always the same.

The IMPORTS section of the mib is sometimes called the linkage section. It lets you import definitions of datatypes and oids from other mibs. Here we get the definition of:

 • mgmt
 • NetworkAddress
 • IpAddress

Nick Urbanik <nicku(at)nicku.org>  ver. 1.1

The Structure of Management Information (SMI) Systems and Network Management

4

 • Counter
 • Gauge
 • TimeTicks

from RFC1155-SMI, the mib from the rfc that defines smiv1.

It also imports OBJECT-TYPE from RFC-1212, the Concise MIB Definition, which defines how mib files are written.

mib-2      OBJECT IDENTIFIER ::= { mgmt 1 }

The line above says that the oid of mib-2 is 1.3.6.1.2.1. RFC1155-SMI defines mgmt as the oid 1.3.6.1.2.

-- groups in MIB-II
system       OBJECT IDENTIFIER ::= { mib-2 1 }
interfaces   OBJECT IDENTIFIER ::= { mib-2 2 }
at           OBJECT IDENTIFIER ::= { mib-2 3 }
ip           OBJECT IDENTIFIER ::= { mib-2 4 }
icmp         OBJECT IDENTIFIER ::= { mib-2 5 }
tcp          OBJECT IDENTIFIER ::= { mib-2 6 }
udp          OBJECT IDENTIFIER ::= { mib-2 7 }
egp          OBJECT IDENTIFIER ::= { mib-2 8 }
transmission OBJECT IDENTIFIER ::= { mib-2 10 }
snmp         OBJECT IDENTIFIER ::= { mib-2 11 }

So here the system group is defines as the oid 1.3.6.1.2.1.1, and so on. A comment is a line starting with --.

-- the Interfaces table

The Structure of Management Information (SMI) Systems and Network Management

5

-- The Interfaces table contains information on the entity's
-- interfaces.  Each interface is thought of as being
-- attached to a `subnetwork'.  Note that this term should
-- not be confused with `subnet' which refers to an
-- addressing partitioning scheme used in the Internet suite
-- of protocols.
ifTable OBJECT-TYPE
    SYNTAX  SEQUENCE OF IfEntry
    ACCESS  not-accessible
    STATUS  mandatory
    DESCRIPTION
            "A list of interface entries.  The number of
            entries is given by the value of ifNumber."
    ::= { interfaces 2 }

This is the first managed object shown here. ifTable represents a table of network interfaces on a managed device. Notice that object names are defined with mixed case, the first letter is lowercase.

Notice that this follows the format of an OBJECT-TYPE in section 1.6 on page 2.

  The SYNTAX of ifTable is SEQUENCE OF IfEntry.

The object is not-accessible, which means that you cannot query the agent for the value of this object. It has a STATUS of mandatory, which means that if an agent complies wiht the mibb-ii specification, then it must implement this object. The DESCRIPTION tells you what this object is. The unique oid is 1.3.6.1.2.1.2.2,

Nick Urbanik <nicku(at)nicku.org>  ver. 1.1

The Structure of Management Information (SMI) Systems and Network Management

6

or iso.org.dod.internet.mgmnt.interfaces.2.

Next, let’s look at the SEQUENCE definition, which is used with the SEQUENCE OF type in the ifTable definition.

IfEntry ::=
SEQUENCE {
ifIndex
    INTEGER,
ifDescr
    DisplayString,
ifType
    INTEGER,
ifMtu
    INTEGER,
ifSpeed
    Gauge,
ifPhysAddress
    PhysAddress,
ifAdminStatus
    INTEGER,
ifOperStatus
    INTEGER,
ifLastChange
    TimeTicks,
ifInOctets
    Counter,
ifInUcastPkts
    Counter,
ifInNUcastPkts

Nick Urbanik <nicku(at)nicku.org>  ver. 1.1

The Structure of Management Information (SMI) Systems and Network Management

7

Counter,
ifInDiscards
Counter,
ifInErrors
Counter,
ifInUnknownProtos
Counter,
ifOutOctets
Counter,
ifOutUcastPkts
Counter,
ifOutNUcastPkts
Counter,
ifOutDiscards
Counter,
ifOutErrors
Counter,
ifOutQLen
Gauge,
ifSpecific
OBJECT IDENTIFIER
}

The name of the SEQUENCE (IfEntry) is mixed-case, but the first letter is capitalised, which is different from the object definition for ifTable. A SEQUENCE is a list of objects that go into one row of a table. After this, we must have OBJECT-TYPE definitions that define each of these variables. A table can have any number of rows. The agent manages the number of rows. An nms can

Nick Urbanik <nicku(at)nicku.org>  ver. 1.1

The Structure of Management Information (SMI) Systems and Network Management

8

also add rows to a table using a set operation.

IfEntry is the data type; rather like a struct definition in the C language.

Let’s look at ifEntry, the definition of what we find in the table, the actual rows of the table themselves. It looks almost the same as the definition for ifTable, except that it has a new clause, INDEX. The index is a unique value that identifies a single row in the table, like an array index. A table is rather like an array of structs in C. The agent assigns these index values. If a router has eight interfaces, then ifTable will contain eight rows.

ifEntry OBJECT-TYPE
    SYNTAX  IfEntry
    ACCESS  not-accessible
    STATUS  mandatory
    DESCRIPTION
            "An interface entry containing objects at the
            subnetwork layer and below for a particular
            interface."
    INDEX   { ifIndex }
    ::= { ifTable 1 }

Here we now look at the definition for ifIndex, the first item in IfEntry. Notice that indexes start from 1.

ifIndex OBJECT-TYPE
    SYNTAX  INTEGER
    ACCESS  read-only

Nick Urbanik <nicku(at)nicku.org>  ver. 1.1

The Structure of Management Information (SMI) Systems and Network Management

9

STATUS  mandatory
DESCRIPTION
"A unique value for each interface.  Its value
ranges between 1 and the value of ifNumber.  The
value for each interface must remain constant at
least from one re-initialization of the entity's
network management system to the next re-
initialization."
::= { ifEntry 1 }

This object is read-only, which means that you can see the value, but not change it.

Here is the last object we look at from this table:

ifDescr OBJECT-TYPE
SYNTAX  DisplayString (SIZE (0..255))
ACCESS  read-only
STATUS  mandatory
DESCRIPTION
"A textual string containing information about the
interface.  This string should include the name of
the manufacturer, the product name and the version
of the hardware interface."
::= { ifEntry 2 }
END

ifDescr is just a textual description of the interface. The mib definition finishes with END.

The Structure of Management Information (SMI) Systems and Network Management

10

2.1 Where can I get the standards documents from?

The standard for smiv1 can be downloaded from

ftp://ftp.rfc-editor.org/in-notes/rfc1155.txt,

and for —SMIv2 at

ftp://ftp.rfc-editor.org/in-notes/rfc2578.txt.

The standards for asn.1 and ber can be downloaded

from http://asn1.elibel.tm.fr/en/standards/.